Find the indicated probability
If P(A or B) = 0.7, P(A) = 0.3, and P(A and B) = 0.15, find P(B)
Thank you

Answer by tommyt3rd(5050)

P(A or B) = P(A) + P(B) - P(A and B)
P(B) = P(A and B)+ P(A or B)-P(A)=
0.15 + 0.70 - 0.3 =
0.55


:)